HUBBLE SPACE TELESCOPE DAILY REPORT #1542 PERIOD COVERED: 0000Z (UTC) 12/21/95 - 0000Z (UTC) 12/22/95 Daily Status Report as of 356/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 6339 (Verification of the UV Spectrum of the HST = Background Light ) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet observations of the HST background light. These FOS = observations were taken parallel to Hubble Deep field (HDF) survey. Both = the HDF and these FOS data will be used to test and improve the existing = engineering model of the HST background light. The observations were = completed as planned, and no anomalies were reported. 1.2 Completed WF/PC-2 5972 (Determination of the Extragalactic Distance = Scale) The WF/PC-2 was used to make observations of the galaxy NGC-3198 = looking for standard candle Cepheid variable stars. The purpose of this = observations is to help estimate the Hubble constant with an accuracy of = 10%. The observations were executed as scheduled, and no anomalies were = noted. 1.3 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration = images. Both photometric and narrow-band filters were used. The = observations were executed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6337 (The Hubble Deep Field Observations) The WF/PC-2 were used to take the Hubble Deep Field (HDF) images. = This project consists of hundreds of long WF/PC-2 exposures of the high = galactic latitude field (HDF-123646), taken in eight visits. A dark = calibration image was also taken. The observations were completed as = planned, and no problems were reported. 1.5 Completed FOS/RD 6342 (The Spectrum of Background Light) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of back ground light. These observations were = taken in parallel during portions of the Hubble Deep Field (HDF) survey. = The data will be used to measure the extragalactic background light (EBL). = The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 1 Successful : 1 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 1 Successful : 1 2.3 Operations Notes: A quite day, no operational problems to report. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Over the upcoming Christmas through New Year's Holidays this Daily = Report will be produced on Dec. 27 (Wednesday) and Dec. 28 (Thursday) = only. Regular coverage will pick up again of Jan. 02, 1996. /DMH